40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 30 18:06:22 2022, Paco, Update, ALS, Optimum ALS recovery - part I als_ool_06_2022.png
    Reply  Thu Jun 30 19:04:55 2022, rana, Summary, PSL, PSL HEPA: How what when why 
Message ID: 16965     Entry time: Thu Jun 30 18:06:22 2022     Reply to this: 16966
Author: Paco 
Type: Update 
Category: ALS 
Subject: Optimum ALS recovery - part I 

[Paco]

In the morning I took some time to align the AUX beams in the XEND table. Later in the afternoon, I did the same on the YEND table. I then locked the AUX beams to the arm cavities while they were stabilized using POX/POY and turned off the PSL hepa off temporarily (this should be turned on after today's work).

After checking the the temperature slider sign on the spectrum analyzer of the control room I took some out-of-loop measurements of both ALS beatnotes (Attachment #1) by running diaggui /users/Templates/ALS/ALS_outOfLoop_Ref_DQ.xml and by comparing them against their old references (red vs magenta and blue vs cyan); it seems that YAUX is not doing too bad, but XAUX has increased residual noise around and above 100 Hz; perhaps as a result of the ongoing ALS SURF loop investigations? It does look like the OLTF UGF has dropped by half from ~ 11 kHz to ~ 5.5 kHz.

Anyways let this be a reference measurement for current locking tasks, as well as for ongoing SURF projects.

Attachment 1: als_ool_06_2022.png  45 kB  Uploaded Thu Jun 30 19:15:46 2022  | Hide | Hide all
als_ool_06_2022.png
ELOG V3.1.3-